Tronic Financial corporation is a financial services holding company headquartered in Ithaca,New York, that offers banking insurance and wealth management service. It pays cash dividends quarterly and also issues stock dividends periodically.

1. At March 31, 2012, True had 9,571,200 shares issued with a par value of $4.00 per share and $62,600 share held in treasury. On April 25, 2012, the company announced that its Board of Directors approved payment of a regular quarterly cash dividend of .75 per share , payable on May 15, 2012, to common shareholder of record on May 7,2012. Assume no shares were acquired or sold by the company after March 31. Give the journal entry to record the declaration of the cash dividend.

2. At December 31, 2009, True reported 2,812,300 shares issued with a par value of $4.00 per share and 17,800 share held in treasury. On January 27,2010, the company announced that its board of directors approved payment of a regular quarterly cash dividend of $.85 per share, payable on February 25, 2010 to common shareholders of record on February 5, 2010. The board also approved the payment of 9% stock dividend distributable on February 25, 2010, to common shareholders of record on February 5, 2010. The share price was $22 when the stock dividend was issued. Assume no treasury shares were acquired or sold after June 30. Prepare the journal entry to record true 's stock dividend.

3. Tronic issued 9% stock dividends in 1995, 2003, 2005, 2006, 2010. In 1998, True issued a three-for-two split. If an investor purchased 800 shares in 1994, how many shares would the investor have in 2012? Fill in the chart below: Shares at the end of: 1994 2,000 1995 1998 2003 2005 2006 2007 2008 2009 2010 Fill in the blank: If an investor 2,000 share in 1994, the investor would have ____________shares in 2012
Answer #1

ANSWER

Share outstanding = 9571200-62600 = 9508600 Shares

1)

Journal entry

Date account and explanation Debit Credit
Apr 25 Cash dividend (9508600*.75) 7131450
Dividend payable 7131450
(To record dividend payable)

2)

Journal entry

No Account and explanation Debit Credit
a Stock dividend (2812300-17800)*9%*22 5533110
Common Stock dividend distributable 1006020
Paid in Capital in excess of par value-Common Stock 4527090
b Common Stock dividend distributable 1006020
Common Stock 1006020

3)

Number of shares in 1994 = 800

Stock dividend in 1995 = 9%

= 800 x 9%

= 72

Number of shares in 1995 = Number of shares in 1994 + Stock dividend in 1995

= 800 + 72

= 872

In 1998, True issued a three-for-two split.

Thus, in 1998, number of shares became = 872 x 3/2

= 1,308

Stock dividend in 2003 = 9%

= 1,308 x 9%

= 118 (rounded to nearest whole number)

Number of shares in 2003 = Number of shares in 1998 + Stock dividend in 2003

= 1,308 + 118

= 1,426

Stock dividend in 2005 = 9%

= 1,426 x 9%

= 128 (rounded to nearest whole number)

Number of shares in 2005 = Number of shares in 2003 + Stock dividend in 2005

= 1,426 + 128

= 1,554

Stock dividend in 2006 = 9%

= 1,554 x 9%

= 140 (rounded to nearest whole number)

Number of shares in 2006 = Number of shares in 2005 + Stock dividend in 2006

= 1,554 + 140

= 1,694

Stock dividend in 2010 = 9%

= 1,694 x 9%

= 152 (rounded to nearest whole number)

Number of shares in 2010 = Number of shares in 2006 + Stock dividend in 2010

= 1,694 + 152

= 1,846

In 2012, investor would have 1,846 shares.

Exact answer may slightly differ due to rounding off.

---------------------------------------------------------------------

DEAR STUDENT,
If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************